During April, the production department of a process manufacturing system completed a number of units of a product and transferred them to finished goods. Of these transferred units, 60,000 were in process in the production department at the beginning of April and 240,000 were started and completed in April. April's beginning inventory units were 60% complete with respect to materials and 40% complete with respect to conversion. At the end of April, 82,000 additional units were in process in the production department and were 80% complete with respect to materials and 30% complete with respect to conversion. 1. Compute the number of units transferred to finished goods. 2. Compute the number of equivalent units with respect to both materials used and conversion used in the production department for April using the weighted-average method.
Answer #1

1) Number of units transferred out = 60000+240000 = 300000 Units

2) Equivalent unit

EUP-Material EUP-Conversion
Unit transferred out 300000 300000
Ending WIP 82000*80% = 65600 82000*30% = 24600
Equivalent unit of production 365600 324600